using a calculator o approximate two values of theta in radians when (theta is less than equal to zero and less than 2 pie) if sec theta = -2.
Answer by stanbon(75887)

using a calculator o approximate two values of theta in radians when (theta is greater than or equal to zero and less than 2 pie)
if sec theta = -2
---
Comment: Way to think about the problem.
If sec(theta) = -2, cos(theta) = -1/2
If cos(theta) were 1/2, theta would be pi/3
But cos(theta) is -1/2, so theta is pi +- (pi/3) = (2/3)pi or (4/3)pi
---
invcos(-1/2) = (2/3)pi or (4/3)pi
===============
Cheers,
Stan H.
===============